Business and Finance. Kevin earned $165 interest for 1 year on an investment of $1500. At the same rate, what amount of interest would be earned by an investment of $2500?
Answer by checkley75(3666)

165/1500=x/2500 now cross multiply
1500x=165*2500
1500x=412500
x=412500/1500
x= 275 would be the interest of an investment of 2500 with the same rate.